# Remarks on the maximal regularity for parabolic boundary value problems with inhomogeneous data

### v1

@@ 1. Introduction
The related $W^{2,1}_{p,\rho}$ parabolic regularity estimate of $v$ in terms of its boundary data $g$ was first studied by Weidemaier [13, 14] and Denk-Hieber-Prüss [7], for $1<p,\rho<\infty$. For the borderline index $p,\rho=1$, Ogawa-Shimizu [9] and [10, Theorems 2.1 & 2.2] established the following $L^{1}$-maximal regularity estimate for $1<p<\infty$,
@@
$\left\|\partial_{t}v,\nabla^{2}v\right\|_{L^{1}(\mathbb{R}_{+};L^{p}(\mathbb{R}^{d}_{+}))}\lesssim\left\|g\right\|_{\accentset{\mbox{\large.}}{F}^{1-\frac{1}{2p}}_{1,1}(\mathbb{R}_{+};L^{p}(\mathbb{R}^{d-1}))}+\left\|g\right\|_{L^{1}(\mathbb{R}_{+};\accentset{\mbox{\large.}}{B}^{2-\frac{1}{p}}_{p,1}(\mathbb{R}^{d-1}))},$ (1.6)
where $\accentset{\mbox{\large.}}{F}^{s}_{p,q}$ is the Lizorkin-Triebel space. Inequality (1.6) is similar to our regularity estimate (1.3) when $m=2$ and $p=1$, and is different from (1.3) in that (1.6) has a mixed exponent. Ogawa-Shimizu also established a trace estimate in [10, Theorem 7.1]. These results can be applied to prove the global well-posedness of free boundary problems for Navier-Stokes equations [11].
@@
A closer inspection of the proofs of [6, Lemma 2.8] and [10, Theorems 2.2 & 7.1] reveals that they share the same spirit. We were thus motivated to find an estimate that encloses both (1.6) and (1.7), but it has not been achieved yet. In this aspect, it is desired that $p=1$ is allowed in both (1.3) and (1.4).
@@ 2. Function spaces and interpolation
We first recall definitions and properties of classical homogeneous Sobolev, Bessel potential and Besov spaces in §2.1. We then recall their anisotropic versions in §2.2.
@@ 2.1. Isotropic case
See [12, 2.2.2, 2.3.1, 2.3.5, 3.4.2] for related results concerning the inhomogeneous case. It’s easy to see (i) is true for domain $\mathbb{R}^{d}$, and hence is true for $\Omega$. For (ii) and (iii), it’s easy to see it’s true for domain $\mathbb{R}^{d}$, and by high-order extension from $\mathbb{R}^{d}_{+}$ to $\mathbb{R}^{d}$ (which keeps the $m$-th derivative on the boundary $\partial\mathbb{R}^{d}_{+}$), it’s also true in $\mathbb{R}^{d}_{+}$. ∎
@@
The domain $\mathbb{R}^{d}_{+}$ in Lemma 2.2 (ii) cannot be replaced by a bounded domain $\Omega$, since $u=\hbox{const}\neq 0$ has nontrivial $\accentset{\mbox{\large.}}{W}^{m}_{p}(\Omega)$ norm. It’s similar for Lemma 2.2 (iii). See [12, Proposition 3.4.2] for the explicit nonhomogeneous norm $B^{s}_{p,p}(\Omega)$ in bounded domain. See [2, Theorem 2.37] for the corresponding form of (2.1) when $s$ is an integer.
@@
(i),(ii) and (iii) follow directly from the inhomogeneous version in [12, §3.3.6] (note $\accentset{\mbox{\large.}}{H}^{s}_{p}=\accentset{\mbox{\large.}}{F}^{s}_{p,2}$ in [12, §2.5.6]). See [3, §6.3, 6.4] for $p=1,\infty$ cases of $(iii)_{2}$. (iv) follows from [8, Theorem 17.30]. ∎
Bessel potential space $\accentset{\mbox{\large.}}{H}^{s}_{p}$ and Besov space $\accentset{\mbox{\large.}}{B}^{s}_{p,p}$ are two ways to quantify the properties of the fractional derivatives. Sometimes they are used to define the so-call fractional Sobolev space. The advantage of Bessel potential space is that it matches with the usual Sobolev space when $s$ is an integer. The advantage of Besov space is that it has an explicit form (2.1) of the norm (and [12, Proposition 3.4.2] for $B^{s}_{p,p}$) when $s$ is a non-integer real number.

### v2

@@ Abstract.
Key words: Parabolic boundary value problems, maximal regularity, anisotropic Besov spaces
AMS Subject Classification (2020): 35K20, 46E35, 46B25
@@ 1. Introduction
For corresponding estimates in the time-independent case, see Mironescu-Russ [9, Theorem 1.9], and Leoni [8, Theorems 18.53, 18.57]. The related $W^{2,1}_{p,\rho}$ parabolic regularity estimate of $v$ in terms of its boundary data $g$ was first studied by Weidemaier [14, 15] and Denk-Hieber-Prüss [7], for $1<p,\rho<\infty$. For the borderline index $p,\rho=1$, Ogawa-Shimizu [10] and [11, Theorems 2.1 & 2.2] established the following $L^{1}$-maximal regularity estimate for $1<p<\infty$,
@@
$\left\|\partial_{t}v,\nabla^{2}v\right\|_{L^{1}(\mathbb{R}_{+};L^{p}(\mathbb{R}^{d}_{+}))}\lesssim\left\|g\right\|_{\accentset{\mbox{\large.}}{F}^{1-\frac{1}{2p}}_{1,1}(\mathbb{R}_{+};L^{p}(\mathbb{R}^{d-1}))}+\left\|g\right\|_{L^{1}(\mathbb{R}_{+};B^{2-\frac{1}{p}}_{p,1}(\mathbb{R}^{d-1}))},$ (1.6)
where $\accentset{\mbox{\large.}}{F}^{s}_{p,q}$ is the Lizorkin-Triebel space. Inequality (1.6) is similar to our regularity estimate (1.3) when $m=2$ and $p=1$, and is different from (1.3) in that (1.6) has a mixed exponent. Ogawa-Shimizu also established a trace estimate in [11, Theorem 7.1]. These results can be applied to prove the global well-posedness of free boundary problems for Navier-Stokes equations [12].
@@
A closer inspection of the proofs of [6, Lemma 2.8] and [11, Theorems 2.2 & 7.1] reveals that they share the same spirit. We were thus motivated to find an estimate that encloses both (1.6) and (1.7), but it has not been achieved yet. In this aspect, it is desired that $p=1$ is allowed in both (1.3) and (1.4).
@@ 2. Function spaces and interpolation
We first recall definitions and properties of classical homogeneous isotropic Sobolev, Bessel potential and Besov spaces in §2.1. We then recall their anisotropic versions in §2.2.
@@ 2.1. Isotropic case
See [13, 2.2.2, 2.3.1, 2.3.5, 3.4.2] for related results concerning the inhomogeneous case. It’s easy to see (i) is true for domain $\mathbb{R}^{d}$, and hence is true for $\Omega$. For (ii) and (iii), it’s easy to see it’s true for domain $\mathbb{R}^{d}$, and by high-order extension from $\mathbb{R}^{d}_{+}$ to $\mathbb{R}^{d}$ (which keeps the $m$-th derivative on the boundary $\partial\mathbb{R}^{d}_{+}$), it’s also true in $\mathbb{R}^{d}_{+}$. ∎
@@
The domain $\mathbb{R}^{d}_{+}$ in Lemma 2.2 (ii) cannot be replaced by a bounded domain $\Omega$, since $u=\hbox{const}\neq 0$ has nontrivial $\accentset{\mbox{\large.}}{W}^{m}_{p}(\Omega)$ norm. It’s similar for Lemma 2.2 (iii). See [13, Proposition 3.4.2] for the explicit nonhomogeneous norm $B^{s}_{p,p}(\Omega)$ in bounded domain. See [2, Theorem 2.37] for the corresponding form of (2.1) when $s$ is an integer.
@@
(i),(ii) and (iii) follow directly from the inhomogeneous version in [13, §3.3.6] (note $\accentset{\mbox{\large.}}{H}^{s}_{p}=\accentset{\mbox{\large.}}{F}^{s}_{p,2}$ in [13, §2.5.6]). See [3, §6.3, 6.4] for $p=1,\infty$ cases of $(iii)_{2}$. (iv) follows from [8, Theorem 17.30]. ∎
Bessel potential space $\accentset{\mbox{\large.}}{H}^{s}_{p}$ and Besov space $\accentset{\mbox{\large.}}{B}^{s}_{p,p}$ are two ways to quantify the properties of the fractional derivatives. Sometimes they are used to define the so-call fractional Sobolev space. The advantage of Bessel potential space is that it matches with the usual Sobolev space when $s$ is an integer. The advantage of Besov space is that it has an explicit form (2.1) of the norm (and [13, Proposition 3.4.2] for $B^{s}_{p,p}$) when $s$ is a non-integer real number.